2. You paid $99,500 for a $100,000 T-bill maturing in 60 days. If you hold it until maturity, what is the T-bill yield? What is the T-bill discount?
Answer #1

Face Value FV = $100000

Purchase Price PP = 99500

Number of days to maturity n = 60

Hence, Yield = (FV-PP)/PP * (360/n) = (100000 - 99500)/99500 * (360/60) *100% = 3.015 %

Discount % = (FV-PP)/FV * (360/n) = (100000 - 99500)/100000 * (360/60) *100% = 3 %